OpenStudy (anonymous):

help pweeeeeeeeeeeeeease In 1994, 24 samples from the archeological site at Oslonki, Poland were dated using carbon 14. The archeological site is significant for the large quantities of copper implements and jewelry found there. One of the samples had 47.8% of the carbon 14 that would normally be present in a living organism. Approximately, how old is the sample? (Carbon 14 decays approximately 1.202% every 100 years.)

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

1.202%?

OpenStudy (anonymous):

8319.47

OpenStudy (anonymous):

=(

OpenStudy (anonymous):

im wondering if theres a way to use logs for this

OpenStudy (australopithecus):

One second let me look up the formula for compound decay

OpenStudy (anonymous):

kk thank u

OpenStudy (australopithecus):

Use this

OpenStudy (anonymous):

wait so are we using the slope formula for this

OpenStudy (anonymous):

ok

OpenStudy (australopithecus):

so we have y = m_0e^(rx) where x = time y = amount at time t m_0 = initial amount of carbon 14 which would be 100% r = rate of decay we know m_0 = 100% r = ? we know point (100, 1.202%) so we can solve for r

OpenStudy (australopithecus):

once you solve for r you can simply sub the percentage into y and solve for x to find the time it took to decay

OpenStudy (australopithecus):

follow?

OpenStudy (anonymous):

hmmmmm well the thing is i never used that formula u gave me so im tryna let it sink in

OpenStudy (anonymous):

we used a different method in class we used excel =/

OpenStudy (australopithecus):

so you never saw it in class? because this would exponential decay since it is radioactive

OpenStudy (anonymous):

no we used excel

OpenStudy (australopithecus):

Well this formula is pretty common in grade 12 functions so I dont know and besides the algebra isn't hard it cant hurt to learn it now

OpenStudy (anonymous):

ok gotcha

OpenStudy (australopithecus):

Make an attempt at solving for r and I will tell you if you are on the right track, just use logarithms to solve

OpenStudy (anonymous):

yes we used logs im trying to refresh my memory how would i go about using logs that is way more simpler

OpenStudy (australopithecus):

first divide both sides by 100 then take the natural logarithm of both sides

OpenStudy (australopithecus):

natural logarithm being ln(x)

OpenStudy (anonymous):

log100= log 1.01202

OpenStudy (australopithecus):

remember the rule ln(e^x) = x

OpenStudy (australopithecus):

log100=/= log 1.01202

OpenStudy (anonymous):

so its 385.42

OpenStudy (anonymous):

the answer is 6100 years though i dont know where i messed up

OpenStudy (australopithecus):

sorry I made a mistake the point we have is (100, 98.798%)

OpenStudy (anonymous):

where did u get that number from

OpenStudy (australopithecus):

We also know that at 100 years we have 100%-1.202% =98.798% sample left so thus we have another point (100, 98.798%)

OpenStudy (anonymous):

oh ok and im guessing since its decay your subtracting?

OpenStudy (australopithecus):

after the first initial 100 years we have 1.202% of the sample gone thus we are left with 98.798% of the sample left ok I got it

OpenStudy (anonymous):

It's looking for an approximation and gives the decay rate with a precision of ±100 years. I think \(\large 0.478=1.000(1-0.01202)^t\) with t in units of hundreds of years ought to do it.

OpenStudy (australopithecus):

so you know at t = 0 y = 100% so we have that m_0 = 100% so we have the equation is \[y = 100e^{-rt}\] where 100 = 100% now we sub in the point we have into the equation to solve for the rate or r since we know the point y = 100% - 1.202% = 98.798% x = 100 years we can sub that point into the equation to solve for r \[98.798 = 100e^{-100r}\] You should get r = 0.000120928 Then we have the final formula \[y = 100e^{-0.000120928(t)}\] with this formula/function we can simply sub in the % carbon 14 of the sample to find its date

OpenStudy (australopithecus):

sorry for taking so long to answer this but wolframalpha was acting really strange

OpenStudy (anonymous):

its fine im just reviewing everything u wrote down

OpenStudy (anonymous):

Try \(\large t=\frac{log(0.478)}{log(0.98798)}\)

OpenStudy (anonymous):

oh ok i guess u cant use logs for this then

OpenStudy (anonymous):

oh nvm i guess we can

OpenStudy (anonymous):

61.04 is what i got

OpenStudy (australopithecus):

sorry that is incorrect yomamabf

OpenStudy (anonymous):

t is in units of 100 years . . .

OpenStudy (australopithecus):

seriously follow my method, not to be a jerk to cliff sledge but I got a plausible answer and this equation is for radioactive decay

OpenStudy (australopithecus):

your answer is correct than cliffsledge

OpenStudy (anonymous):

You're not being a jerk at all, @Australopithecus , I just think you are working to hard on a question that asks for an approximation with a margin of error of ±100 years.

OpenStudy (australopithecus):

your answer is correct then cliffsledge

OpenStudy (australopithecus):

you are probably right cliffsledge ha

OpenStudy (australopithecus):

but it really isn't that hard to use this formula, my laziness just got the better of me

OpenStudy (anonymous):

I think what you're doing is probably correct too. Here is a summary of an alternate method: Generic exponential model, \(\large A=P(1+r)^t\) A=0.478, the 47.8% remaining P=1.000, the 100% started with r=-0.01202, the 1.202% decay rate t= time (in 100s of years). Plug in the knowns and solve for t.

OpenStudy (anonymous):

yeaaaaaaaaaaaa thats what we used in class

OpenStudy (anonymous):

Since it's approximate to ±100 years, a continuous compounding model isn't appropriate.

OpenStudy (anonymous):

If a half-life was given, then using e and natural logs would work nicely.

OpenStudy (anonymous):

the answer however is 6100 years and its not adding up

OpenStudy (anonymous):

Why not? I got 6104 years when I did it.

OpenStudy (anonymous):

hang on lemme try this again

OpenStudy (anonymous):

sorry i plugged everything in and i got a .47232

OpenStudy (anonymous):

? Hmm.. \(\large A=P(1+r)^t\) \(\large 0.478=(1-0.01202)^t\) \(\large log_{0.98798}(0.478)=log_{0.98798}((1-0.01202)^t)\) \(\large t= log_{0.98798}(0.478) = log(0.478) \div log(0.98798)\)

OpenStudy (anonymous):

oh u used logs? ok ill try it again

OpenStudy (anonymous):

Have to. Need to solve for the exponent, that's what logs do.

OpenStudy (anonymous):

ohhhh i finally got it thank u so much <333 id give u a medal but nincompoop stole it and deleted his replies

OpenStudy (anonymous):

You're welcome.